OpenStudy (anonymous):

can someone please help me solve 15x^4-14x^2+3=0

OpenStudy (anonymous):

15x4-9x2-5x2 +3=0 3x2(5x2-3)-(5x2-3)=0

OpenStudy (anonymous):

(3x2-1)(5x2-3)=0

OpenStudy (anonymous):

(3x^2-1)(5x^2-3)=0 +-Sqrt[3/5] +-1/Sqrt[3]

OpenStudy (anonymous):

u r cute

OpenStudy (anonymous):

Thanks and how am i suppose to write that answer

OpenStudy (anonymous):

\[x \in\left\{\pm \sqrt{3/5}; \pm1/\sqrt{3} \right\}\]

OpenStudy (anonymous):

15x4-9x2-5x2 +3=0 3x2(5x2-3)-(5x2-3)=0 (3x2-1)(5x2-3)=0 3x2-1=0 , 5x2-3=0 x2=1/3, x2=3/4 x=+-1/sqrt3 , x=+- sqrt3 /2

OpenStudy (anonymous):

where r u from?

OpenStudy (anonymous):

sc

OpenStudy (anonymous):

sc?

OpenStudy (anonymous):

Scotland?

OpenStudy (anonymous):

and u are from

OpenStudy (anonymous):

south cakcalaki

OpenStudy (anonymous):

australia

OpenStudy (anonymous):

whoops :) Im from Czech republic :)

OpenStudy (anonymous):

what grade r u in?

OpenStudy (anonymous):

me or dollar?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

I dont know your educational system. Im on hight school and im finishing it on 2012

OpenStudy (anonymous):

ankur u mean -1 sq root 3, and -3/2

OpenStudy (anonymous):

college

OpenStudy (anonymous):

x=1/sqrt3, +-1/sqrt3 , x= sqrt3 /2, - sqrt3 /2

OpenStudy (anonymous):

you have error there

OpenStudy (anonymous):

it is sqrt(3/5) not (3/4)

OpenStudy (anonymous):

serychj im confused on which answer to put

OpenStudy (anonymous):

mine is correct :) his have error in one line before end: 5x2-3=0 x2=3/4 -> error x^2=3/5

OpenStudy (anonymous):

sorry guys but all the answers were wrong after a couple of tries it generates the correct answer and it was -sq3/3,sq3/3,and -sq15/5,sq15/5 thanks anyway

OpenStudy (anonymous):

check this: http://www.wolframalpha.com/input/?i=15x^4-14x^2%2B3%3D0 your wrong